# On the Weight Spectrum of Rate-Compatible Polar Codes

### v2

@@ IV-C Number of minimum-weight codewords in Wang-Liu shortened polar codes
For Wang-Liu shortening, a codeword $\bm{c}$ in $T(f)$ is not shortened if and only if $\bm{c}_{[2^{m}-i,2^{m}-1]}=0$. In other words, it must have weight $2^{m-r}$ after the last $a$ bits punctured.
The theorem is a direct consequence of the symmetry: puncturing the last $i$ bits yields the identical weight spectrum as puncturing the first $i$ bits through the affine transformation $x_{t}\to x_{t}+1$ for all $1\leq t\leq m$ ∎
@@
Suppose $C(\mathcal{I},Y_{i})$ is a shortened decreasing $r$-th order polar code with shortening pattern $Y_{i}$ and minimum distance $d=2^{m-r}$, where $Y_{i}$ is the last $i$ bits of the QU sequence $\bm{q}$. Then $A_{d}(C(\mathcal{I},Y_{i}))=\sum_{f\in\mathcal{I}_{r}}P_{f}(d,i)$.
@@
Let $C(\mathcal{I},Y_{2})$ be the code with the length $N=8$, the information set =$\mathcal{I}=\{x_{1}x_{2},x_{2}\}$ and last two bits shortened. Then $d_{\min}(C(\mathcal{I},Y_{2}))=2$. From Theorem 4 and 5, $A_{2}(C(\mathcal{I},Y_{2}))=P_{x_{1}x_{2}}(2,2)=2$, where $P_{x_{1}x_{2}}(2,2)$ can be derived by Algorithm 2.
@@ V Average weight spectrum of pre-transformed rate-compatible polar codes
In this section, we introduce efficient formulas for calculating the average weight spectrum of rate-matching pre-transformed polar codes. A distinguishing attribute of our approach is its versatility-it operates without the limitations imposed by partial ordering or specific rate matching patterns. Furthermore, our approach enables a calculation of the complete weight spectrum. The key idea of our approach lies in mitigating the complexity of the enumeration problem through averaging over a appropriate chosen ensemble of codes. Specifically, we derive the average weight spectrum by taking the expectation over a set of random pre-transformation matrices with $i.i.d.$ upper triangular components. This strategy significantly reduce the number of unique terms to a linear scale, thus circumventing the exponential computational complexity typically associated with the calculation of specific weight spectrum.
To begin our analysis, we provide a recursive computation of the weight spectrum for polar cosets. For clarity, we dissect the coset into its constituent odd and even parts. This dissection allows us to decompose the length $N$ code as a direct sum of two codes [37, §$\S 9$ of ch.2], each of half the original length, i.e., $\frac{N}{2}$. This approach is similar to the technique employed in [26] for calculating the spectrum of coset codes, albeit with a focus on the mother code length. Here, we extend this concept to the case under rate-matching.
@@
$=\sum_{d_{1}+d_{2}=d}A_{d_{1}}(m,(\bm{u}_{1,e}^{2i-2},0),X_{\hat{e}})A_{d_{2}}(m,(\bm{u}_{1,o}^{2i-2}\oplus\bm{u}_{1,e}^{2i-2},u_{2i-1}),X_{\hat{o}})$
$+\sum_{d_{1}+d_{2}=d}A_{d_{1}}(m,(\bm{u}_{1,e}^{2i-2},1),X_{\hat{e}})A_{d_{2}}(m(\bm{u}_{1,o}^{2i-2}\oplus\bm{u}_{1,e}^{2i-2},u_{2i-1}+1),X_{\hat{o}})$ (12)
The recursive formulas for the weight spectrum of punctured polar cosets are presented in equations (11) and (6). The shortening case follows the same recursive manners by replacing the puncturing pattern $X$ by the shortening pattern $Y$. The only difference is the boundary conditions when $m=0$, i.e., the code length is equal to 1. If the code bit is neither punctured nor shortened,
@@
Since the proof of Theorem 6 is similar as [26, Theorem 1], we only illustrate the key points of the proof.
@@
Drawing from the proof in [26, Theorem 1], it is established that the odd and even components of a coset code are independent, i.e., $(\bm{u}\bm{F}_{N})_{o}=(\bm{u}_{o}+\bm{u}_{e})\bm{F}_{N/2}=\bm{u}^{\prime}\bm{F}_{N/2}$ and $(\bm{u}\bm{F}_{N})_{e}=\bm{u}_{e}\bm{F}_{N/2}$, where $\bm{u}^{\prime}$ and $\bm{u}_{e}$ are independent with each other. This independence implies that the weight spectrum of a coset of length-$N$ can be determined through the spectrum of two constituent cosets, each of length-$\frac{N}{2}$. In comparison to the mother code length case, the distinctive feature lies in the application of rate-matching patterns to the half-length cosets. Specifically, for the two cosets of length-$\frac{N}{2}$, their respective rate-matching patterns correspond to the odd and even components extracted from the original length-$N$ coset.
As for the boundary condition, it is straightforward for the case without rate-matching. For puncturing case, $\mathcal{C}_{P}(0,\bm{u}_{1},X=\{1\})$ is the coset with one empty codeword, so $A_{0}(0,\bm{u}_{1},X=\{1\})=1$ and $A_{1}(0,\bm{u}_{1},X=\{1\})=0$. For shortening case, $\bm{u}_{1}=1$ contradicts with the shorten pattern $Y=1$, then $\mathcal{C}_{S}(0,\bm{u}_{1}=1,Y=\{1\})$ is empty, so $A_{0}(0,\bm{u}_{1}=1,Y=\{1\})=A_{1}(0,\bm{u}_{1}=1,Y=\{1\})=0$. $\mathcal{C}_{S}(0,\bm{u}_{1}=1,Y=\{1\})$ contains one empty codeword, so $A_{0}(0,\bm{u}_{1}=0,Y=\{1\})=1,A_{1}(0,\bm{u}_{1}=0,Y=\{1\})=0$.
@@
Suppose $m=3$, puncturing pattern $X=\{1,2,5\}$, consider the weight spectrum of coset $\mathcal{C}_{P}(3,(0,0,1,0),X)$. From the proof of Theorem 6, we decompose it into odd and even parts, that is, $\mathcal{C}_{P}(3,(0,0,1,1),X)=\{(\bm{c}_{o},\bm{c}_{e})|\bm{c}_{o}\in\mathcal{C}_{P}(2,(0,1)\oplus(0,0),X_{\hat{o}}=\{1,3\}),\bm{c}_{e}\in\mathcal{C}_{P}(2,(0,0),X_{\hat{e}}=\{1\})\}$. Then $A_{d}(3,(0,0,1,0),X)=\sum_{d_{1}+d_{2}=d}A_{d}(2,(0,1),X_{\hat{o}})A_{d}(2,(0,1),X_{\hat{e}})$.
@@
Let =$\mathcal{I}=\{I_{1},...,I_{K}\}$ be the information set, $I_{j}^{s}=|\{i|Y_{i}>I_{j}\}|$ be the number of shorten bits whose index is large than $I_{j}$. Then
@@
Let ${\bm{f}}_{N}^{(i)}$ be the $i$-$th$ row vector of $\bm{F}_{N}$ and ${\bm{g}}_{N}^{(i)}$ be the $i$-$th$ row vector of $\bm{G}_{N}=\bm{T}\bm{F}_{N}$ after puncturing, where $\bm{T}$ is the random pre-transformation matrix. The proof is similar to [28, Lemma 1]. The only difference is for punctured codes,
@@
The algorithm to calculate the weight spectrum of punctured pre-transformed polar cosets is described in Algorithm 3. When considering the scenario of shortening, the underlying algorithm remains identical. The primary divergence between the two cases lies in the boundary conditions.
@@
0:  the punctured polar code $C_{P}(\mathcal{I},X)$ with code length $N=2^{m}$, dimension $K$ and information set =$\mathcal{I}=\{I_{1},\dots,I_{K}\}$.
@@ VI Numerical Results
Table 2 presents a comparison between $\sum_{f\in\mathcal{I}}P_{f}(w,X)$ calculated in Theorem 3 and the weight spectrum collected through SCL decoding with a large list [18] for weights less than or equal to $2^{m-r}$. The results demonstrate that the lower bound in Theorem 3 could be accurate when the number of punctured bits is small. However, when the number of punctured bits is large, the lower bound might be significantly smaller than the actual value. The reason is that more bits punctured can result in a greater number of codewords with weights exceeding $2^{m-r}$ being rendered with weight less than or equal to $2^{m-r}$.
@@
[…]
@@
Figure 4 and 5 compare the union bounds [14] calculated by the average weight spectrum obtained through our methods with the performance of SCL decoding with list size 32 for rate-matching PC-polar codes. The parity check is implemented by a 5-length cyclic shift register [23], and all frozen bits are used for parity check. The information set is chosen according to 5-$th$ generation New Radio (5G-NR) communications systems [38]. The rate-matching patterns includes QUP, Wang-Liu shortening and sub-block interleaving in 5G-NR communications systems. The simulation results indicate that the union bound derived from our results also closely approximates the performance of pre-transformed polar codes, particularly at high SNR.
@@ VII Conclusion
In this paper, we develop a systematic framework for the enumeration of the minimum-weight codewords in QUP, Wang-Liu shortened, and bit-reversal shortened decreasing polar codes. Furthermore, We compute the average weight spectrum of pre-transformed polar codes, applicable to a more expansive rate-matching configurations. Iterative formulas and algorithms are provided with a polynomial complexity relative to code length. Empirical simulations show the capacity of our proposed techniques to approximate the error-correcting performance of rate-matching polar codes with large list size.

### v3

@@ IV-C Number of minimum-weight codewords in Wang-Liu shortened polar codes
For Wang-Liu shortening, a codeword $\bm{c}$ in $T(f)$ is not shortened if and only if $\bm{c}_{[2^{m}-i,2^{m}-1]}=0$. In other words, it must retain its original weight $2^{m-r}$ after the last $a$ bits punctured.
The theorem is a direct consequence of the symmetry: puncturing the last $i$ bits yields the same weight spectrum as puncturing the first $i$ bits through the affine transformation $x_{t}\to x_{t}+1$ for all $1\leq t\leq m$. ∎
@@
Let $C(\mathcal{I},Y_{i})$ be a shortened decreasing $r$-th order polar code with shortening pattern $Y_{i}$ and minimum distance $d=2^{m-r}$, where $Y_{i}$ is the last $i$ bits of the QU sequence $\bm{q}$. Then $A_{d}(C(\mathcal{I},Y_{i}))=\sum_{f\in\mathcal{I}_{r}}P_{f}(d,i)$.
@@
Let $C(\mathcal{I},Y_{2})$ be the code with a length of $N=8$, the information set $\mathcal{I}=\{x_{1}x_{2},x_{2}\}$ and last two bits shortened. Then $d_{\min}(C(\mathcal{I},Y_{2}))=2$. According to Theorem 4 and 5, $A_{2}(C(\mathcal{I},Y_{2}))=P_{x_{1}x_{2}}(2,2)=2$, where $P_{x_{1}x_{2}}(2,2)$ can be determined using Algorithm 2.
@@ V Average weight spectrum of pre-transformed rate-compatible polar codes
In this section, we introduce efficient formulas for calculating the average weight spectrum of rate-matched pre-transformed polar codes. A distinguishing attribute of our approach is its versatility-it operates without the limitations imposed by partial ordering or specific rate matching patterns. Furthermore, our approach enables the calculation of the complete weight spectrum. The key idea of our approach lies in mitigating the complexity of the enumeration problem through averaging over a suitably chosen ensemble of codes. Specifically, we derive the average weight spectrum by taking the expectation over a set of random pre-transformation matrices with $i.i.d.$ upper triangular components. This strategy significantly reduces the number of unique terms to a linear scale, thus circumventing the exponential computational complexity typically associated with calculating specific weight spectrum.
To begin our analysis, we provide a recursive computation of the weight spectrum for polar cosets. For clarity, we break down the coset into its odd and even components. This allows for a decomposition for the length $N$ code as a direct sum of two codes [39, $\S 9$ of ch.2], each of half the original length $\frac{N}{2}$. This approach is similar to the technique employed in [27] for calculating the spectrum of coset codes, albeit with a focus on the mother code length. Here, we extend this method to accommodate rate-matching.
@@
[…]
The recursive formulas for the weight spectrum of punctured polar cosets are presented in Equations (11) and (6). The shortening case follows the same recursive procedure but the puncturing pattern $X$ is replaced by the shortening pattern $Y$. The only difference lies in the boundary conditions when $m=0$, i.e., the code length is equal to 1. If the code bit is neither punctured nor shortened,
@@
Since the proof of Theorem 6 is similar as [27, Theorem 1], we will only illustrate the key points of the proof.
@@
Drawing from the proof in [27, Theorem 1], it is established that the odd and even components of a coset code are independent, i.e., $(\bm{u}\bm{F}_{N})_{o}=(\bm{u}_{o}+\bm{u}_{e})\bm{F}_{N/2}=\bm{u}^{\prime}\bm{F}_{N/2}$ and $(\bm{u}\bm{F}_{N})_{e}=\bm{u}_{e}\bm{F}_{N/2}$, where $\bm{u}^{\prime}$ and $\bm{u}_{e}$ are independent of each other. This independence implies that the weight spectrum of a coset of length-$N$ can be determined through the spectrum of two constituent cosets, each of length-$\frac{N}{2}$. In comparison to the mother code length case, the distinctive feature lies in the application of rate-matching patterns to the half-length cosets. Specifically, for the two cosets of length-$\frac{N}{2}$, their respective rate-matching patterns correspond to the odd and even components extracted from the original length-$N$ coset.
As for the boundary condition, it is straightforward for the case without rate-matching. For the puncturing case, $\mathcal{C}_{P}(0,\bm{u}_{1},X=\{1\})$ is the coset with one empty codeword, so $A_{0}(0,\bm{u}_{1},X=\{1\})=1$ and $A_{1}(0,\bm{u}_{1},X=\{1\})=0$. For the shortening case, $\bm{u}_{1}=1$ contradicts with the shortening pattern $Y=1$, then $\mathcal{C}_{S}(0,\bm{u}_{1}=1,Y=\{1\})$ is empty, so $A_{0}(0,\bm{u}_{1}=1,Y=\{1\})=A_{1}(0,\bm{u}_{1}=1,Y=\{1\})=0$. $\mathcal{C}_{S}(0,\bm{u}_{1}=1,Y=\{1\})$ contains one empty codeword, so $A_{0}(0,\bm{u}_{1}=0,Y=\{1\})=1,A_{1}(0,\bm{u}_{1}=0,Y=\{1\})=0$.
@@
[…]
@@
Let $\mathcal{I}=\{I_{1},...,I_{K}\}$ be the information set, $I_{j}^{s}=|\{i|Y_{i}>I_{j}\}|$ be the number of shorten bits with indices larger than $I_{j}$. Then
@@
Let ${\bm{f}}_{N}^{(i)}$ be the $i$-thth row vector of $\bm{F}_{N}$ and ${\bm{g}}_{N}^{(i)}$ be the $i$-thth row vector of $\bm{G}_{N}=\bm{T}\bm{F}_{N}$ after puncturing, where $\bm{T}$ is the random pre-transformation matrix. The proof is similar to [30, Lemma 1]. The only difference is that for punctured codes,
@@
The algorithm to calculate the weight spectrum of punctured pre-transformed polar cosets is described in Algorithm 3. When considering shortening, the underlying algorithm remains identical. The only divergence between the two cases lies in the boundary conditions.
@@
0: the punctured polar code $C_{P}(\mathcal{I},X)$ with code length $N=2^{m}$, dimension $K$ and information set $\mathcal{I}=\{I_{1},\dots,I_{K}\}$.
@@ VI Numerical Results
Table 2 compares the value of $\sum_{f\in\mathcal{I}}P_{f}(w,X)$ calculated in Theorem 3 with the weight spectrum collected through SCL decoding with a large list [18], for weights less than or equal to $2^{m-r}$. The results demonstrate that the lower bound in Theorem 3 can be accurate when the number of punctured bits is small. However, when the number of punctured bits is large, the lower bound might be significantly smaller than the actual value. The reason is that a greater number of bits punctured can cause codewords with weights exceeding $2^{m-r}$ to be rendered with weights less than or equal to $2^{m-r}$.
@@
Figures 1 to 6 compare the union bounds [14] calculated by the weight spectrum obtained through our methods with the performance of SCL decoding with list size 32 for QUP, Wang-Liu shortened and bit-reversal shortened polar codes with different lengths and rates. Here the union bounds are calculated as $\sum_{d\leq 2^{m-r}}A_{d}Q(-\sqrt{d}/\sigma)$ where $\sigma$ is the standard variance of additive white Gaussian noise and $Q(x)=\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}e^{-z^{2}/2}dz$ is the spectrum function of the standard normal distribution. The information set is chosen by the Gaussian Approximation [38] under rate matching. For shortened codes, the numbers of minimum-weight codewords are derived from Theorem 2 and Theorem 5. For $r$-thth order punctured codes, the numbers of codewords with weight no larger than $2^{m-r}$ are approximated based on Theorem 3. The minimum weight $d_{\min}$ and the number of minimum-weight codewords $A_{d_{\min}}$ are presented in Table 3. The simulation results show that the union bound calculated by our methods closely matches the performance of SC-list decoding, particularly at high SNR.
@@
Figure 7 and 8 compare the union bounds [14] calculated by the average weight spectrum obtained through our methods with the performance of SCL decoding with list size 32 for rate-matching PC-polar codes. The parity check is implemented by a 5-length cyclic shift register [23], and all frozen bits are used for parity check. The information set is chosen according to 5-$th$ generation New Radio (5G-NR) communications systems [40]. The rate-matching patterns include QUP, Wang-Liu shortening and sub-block interleaving in 5G-NR communications systems. The simulation results indicate that the union bound derived from our methods also closely approximates the performance of pre-transformed polar codes, particularly at high SNR.
@@ VII Conclusion
In this paper, we develop a systematic framework for the enumeration of the minimum-weight codewords in QUP, Wang-Liu shortened, and bit-reversal shortened decreasing polar codes. Furthermore, we compute the average weight spectrum of pre-transformed polar codes, applicable to more expansive rate-matching configurations. Iterative formulas and algorithms are provided with a polynomial complexity with respect to code length. Empirical simulations show that our proposed techniques are capable of approximating the error-correcting performance of rate-matching polar codes with large list size.